A plane flies 390 miles with the wild and 325 miles against the wind in the same length of time. If the speed of the wind is 10 mph, find the speed of the plane in still air.
Answer by mananth(16949)

let speed of plane in still air be x mph
speed of wind = 10 mph
..
with wind 390 miles
time taken = 390/x+10
...
against wind = 325 miles
time = 325/x-10
..
both these times are same
...
325/(x-10)=390/(x+10)
390(x-10)=325(x+10)
390x-3900 = 325x +3900
390x-325x=7800
65x = 7800
x= 7800/65
x= 120 mph
....
m.ananth@hotmail.ca